AIDA GELINA BRIKEN nToF CRIB ISOLDE CIRCE nTOFCapture DESPEC DTAS EDI_PSA 179Ta CARME StellarModelling DCF K40
  CARME, Page 21 of 33  ELOG logo
ID Date Author Subject
  264   Thu Feb 24 04:17:35 2022 CB, TD (remotely)MIDAS DAQ no longer working

DAQ server crashed around 24:00

Rebooted server but could not could not rescue remotely - could not contact raspberry pi in control of FEE bias.

Entered cave, rebooted raspberry pi. Would not accept remote connections.

Had to manually start DHCP service on carme-gsi using

systemctl start dhcpd

 

Raspi could then be connected.

Started DAQ, TapeServer and Merger as per usual. Tapeserver refuses writing to disk.

Checked merged - very odd timestamp situation. See attached. Timestamp look OK in the GSI WR tab. See attached.

Power cycled server and FEES

Timestamp issue solved. All are now reading 0x16 etc.

 

However tapeserver still does not write to disk. Reasons unclear. Possibly Merger is not merging and not passing data on?

Have to stop data collection until solved.

 

  263   Wed Feb 23 22:29:50 2022 PE, EM, CB (remotely)check - 23:30

Motor position ok

Data rate: 1030kbytes/s

T ok

System wide check ok

plots uploaded

 

DAQ crashed soon afterwards

  262   Wed Feb 23 21:27:29 2022 TD, CNcheck

Motor position ok

Data rate: 1030kbytes/s

T ok

System wide check ok

plots uploaded

 

  261   Wed Feb 23 20:52:56 2022 CN, TDSystem wide checks
Attachments 1-5 aida01-aida05 options

System wide checks

Attachment 6 - changes to options files - differences reported but none observed (attachments 1-5)

Attachment 7 - collect memory info from FEE64 Linux

Attachment 8 - collect FPGA timestamp errors

Attachment 9 - WR decoder status

Attachment 10 - ADC calibration

Attachment 11 - Clock status
  260   Wed Feb 23 20:06:43 2022 TD CNE141 kinematics
Assuming DSSSD is 1000um thick we can expect to observe protons for E_lab=11MeV as follows

14N(d,p)15N(gs)   E_p=5.4MeV
14N(d,p)15N(5.3)  E_p=8.0MeV
  259   Wed Feb 23 19:22:55 2022 CN, TDcheck

 

- Motor checked , OK

SW3

- Merger OK

- Screenshot uploaded

- Data rate : 755 kbytes/s

SW4

- current plot uploaded

SW2

- T OK

- system wide check OK

- Statistics ok

- Spectra ok

SW5

4 plots uploaded

 

  258   Wed Feb 23 18:58:12 2022 TD CNWednesday 23 February 15.00-23.00 shift

Assorted screenshots of the CARME, CRYRING and gas jet target control systems in operation

Attachment 1 - detector motion controls

Attachment 2 - gas jet target pressures (blue proxy for target density, red proxy for gas load to CRYRING)

Attachment 3 - Schottky probe monitor of beam

Attachment 4 - CRYRING controls monitor

Attachment 5 - CRYRING vacuum pressures

Attachment 6 - CRYRING beam position and profile monitor 
  257   Wed Feb 23 16:57:35 2022 CB, CNOptimising motor motion

Optimised motor motion during run R58

Moved 2 mm in (20.3 potentiometer)

29.47 potentiometer

36.5 potentiometer

(started moving pneumatic)

64.36 potentiometer

74.03 potentiomater (servo in by 10 mm)

 

Moved in, collected counts for 5 minutes, moved out, waited for re-injection.

 

20.3 - 400 counts on the Rutherford peak, bins 125-131 on lowEnergyExTotalDSSD0

20.47 - 646

36.5 - 827

64.36 - 3900

74.03 - 10040

 

Based on Rutherford rate ratio, estimated we are 40-50 mm away from interaction point, i.e. beam roughly 1-2 cm on outer half of ring which may be consistent with BPM measurement in YR08


Decided to use 74.03 mm for data collection.

Started data acquisition on R59 at 17:27

Beam 3E7 at injection, 600 second storage, 3 second between cycles

Target pressure 1.3E12 based on 8.21E18 x S1=1.6E-8

  256   Wed Feb 23 10:56:06 2022 Laszlo, ThanassisCurrent status

11:55
Temps OK
System Check OK
White Rabbit fail diff: 102
Write Speed: ~1.2Mb/s
Leakage Current: ~3.47uA
Cryo Temp = 129.2K setpoint (unchanged)
Pressures stable

AVG Pressures:
YR07: 1.99e-10mbar
YR08: 2.6e-9mbar
YR09: 1.03e-9mbar
YR10: 1.7e-10mbar

  255   Wed Feb 23 09:50:19 2022 Laszlo, ThanassisMoving detector in

We are moving the detector in run 56

  254   Wed Feb 23 08:58:30 2022 Laszlo, ThanassisCurrent status

10:00
Temps OK
System Check OK
White Rabbit fail diff: 102
Write Speed: ~1.2Mb/s
Leakage Current: ~3.47uA
Cryo Temp = 129.2K setpoint (unchanged)
Pressures stable

AVG Pressures:
YR07: 1.93e-10mbar
YR08: 2.64e-9mbar
YR09: 1.13e-9mbar
YR10: 1.55e-10mbar

  253   Wed Feb 23 07:58:38 2022 Laszlo, ThanassisCurrent status

09:00
Temps OK
System Check OK
White Rabbit fail diff: 101
Write Speed: ~1.2Mb/s
Leakage Current: ~3.47uA
Cryo Temp = 129.2K setpoint (unchanged)
Pressures stable

AVG Pressures:
YR07: 1.88e-10mbar
YR08: 2.28e-9mbar
YR09: 9.15e-10mbar
YR10: 1.52e-10mbar

  252   Wed Feb 23 06:58:19 2022 Laszlo, ThanassisCurrent status

08:00
Temps OK
System Check OK
White Rabbit fail diff: 99
Write Speed: ~1.0Mb/s
Leakage Current: ~3.47uA
Cryo Temp = 129.2K setpoint (unchanged)
Pressures stable

AVG Pressures:
YR07: 2.06e-10mbar
YR08: 2.71e-9mbar
YR09: 1.06e-9mbar (up from previous check)
YR10: 1.78e-10mbar

  251   Wed Feb 23 06:00:25 2022 Jordan, GeorgeCurrent status

07:00
Pressure Stable
Temps OK
System Check OK
White Rabbit fail diff: 90
Write speed: ~1Mb/s
Leakage Current: 3.47uA
Cryo Temp Setpoint: 129.2K (unchanged)

AVG Pressures:
YR07: 1.66e-10mbar
YR08: 1.93e-9mbar
YR09: 7.63e-10mbar
YR10: 1.26e-10mbar

  250   Wed Feb 23 04:55:48 2022 Jordan, GeorgeCurrent status

05:55
Temps OK
System Check OK
White Rabbit fail diff: 90
Write Speed: ~1.0Mb/s
Leakage Current: ~3.47uA
Cryo Temp = 129.2K setpoint (unchanged)
Pressures stable

AVG Pressures:
YR07: 1.79e-10mbar
YR08: 2.21e-9mbar
YR09: 8.64e-10mbar
YR10: 1.36e-10mbar

  249   Wed Feb 23 03:52:13 2022 Jordan, GeorgeCurrent status

04:52
Pressure seems to be climbing a little but still stable (just peaked at end of check)
Temps OK
Write Speed: ~1.0Mb/s
System Checks OK
White Rabbit fail diff: 90
Leakage Current: 3.47uA
Cryo Temp: 129.2K (unchanged)

AVG Pressures:
YR07: 1.79e-10mbar
YR08: 2.3e-9mbar
YR09: 9.46e-10mbar
YR10: 1.41e10mbar

Avg pressures plots for last hour and last 6 hours attached.

  248   Wed Feb 23 02:52:09 2022 Jordan, GeorgeCurrent status

03:52

Temps OK

System Checks OK

White rabbit diff: 90

Data rate: 1.2Mb/s

Leakage Current: 3.474uA

Cryo temp = 129.2K (unchanged)

Pressures stable

AVG Pressures:
YR07: 1.74e-10mbar
YR08: 2.15e-9mbar
YR09: 8.61e-10mbar
YR10: 1.32e-10mbar

  247   Wed Feb 23 01:56:59 2022 Jordan, GeorgeCurrent status

02:56

Temps OK

write speed ~1.2Mb/s

System check OK

White rabbit diff - 89

Leakage Current ~3.472uA

Pressures stable

Beam intensity ~3.0e7

cryostat setpoint 129.2K (unchanged)

AVG Pressures:
YR09: 8.48e-10mbar

YR10: 1.28e-10mbar

YR07: 1.71e-10mbar

YR08: 2.1e-9mbar

Aida 1 1*L plot attached. Some channels appear missing. Width of rutherford peaks ~280-300 channels.

  246   Wed Feb 23 00:54:17 2022 Jordan, GeorgeCurrent status

02:04

Temps OK

Data Rate - ~1.1Mb/s

System Check OK

White rabbit diff - 83

Leakage Current - ~3.475uA

YR09 AVG: 7.51E-10mbar

YR10AVG: 1.21E-10mbar

YR07AVG: 1.67E-9mbar

YR08AVG: 1.91E-9mbar

  245   Tue Feb 22 21:57:13 2022 George, JordanCurrent status

23:53

Temperatures OK

System checks OK - white rabbit fail diff ~ 80

Pressures stable

Cryo temp = 129.2 K setpoint

Data rate ~1.2 MB/s

Leakage current ~ 3.47 uA

00:53

YR10 avg = 1.29E-10 mbar

YR09 avg = 8.7E-10 mbar

YR04 avg = 1.13E-11 mbar

YR07 avg = 1.76E-10 mbar

YR08 avg = 2.18E-9 mbar

Data rate ~ 1.1 Mb/s

Temperatures OK

system checks OK - white rabbit fail diff ~ 80

Pressure stable, cryostat temp setpoint = 129.2k

Leakage current = 3.47 uA

Beam intensity reads ~ 2.6-3E7

ELOG V3.1.4-unknown